### v1

@@ 7 Conclusions and Future Work
We proposed a combined simulation and real-world benchmark to enable progress on the important problem of Open-Vocabulary Mobile Manipulation. We ran extensive experiments showing promising simulation and real-world results from two baselines: a heuristic baseline based on a state-of-the-art motion planner [2] and a reinforcement learning baseline trained with DDPPO [60]. In the future, we hope to improve the complexity of the problem space, adding more complex natural language and multi-step commands instead of pick-and-place, and provide end-to-end baselines instead of modular policies. Various proposed solutions for open-vocabulary navigation [9, 11, 12] and manipulation of unknown objects [10, 13, 18, 17] suggest possible ways of improving performance.
@@ Appendix A Extended Related Work
Datasets. In addition to the environments referenced in Table 3, offline datasets including robot interactions with scenes have been used widely to train models. These datasets are typically obtained using robots alone (e.g., [76, 77]), by teleoperation (e.g., [78, 79]) or human-robot demonstration (e.g., [80]). Previous work such as [81] aim to collect large-scale datasets while works such as [82] consider scaling across multiple embodiments. [83] take a step further by collecting robot data in unstructured environments. Unlike these works, we do not limit our users to a specific dataset. Instead, we provide a simulator with various scenes that can generate large-scale consistent data for training. Also, note that we test the models in unseen environments, while most of the mentioned works use the same environment for training and testing.
Simulation benchmarks. The embodied AI community has provided various benchmarks in simulation platforms for tasks such as navigation [1, 84, 85, 86, 87], object manipulation [88, 34, 89, 90], instruction following [6, 91, 92, 93], room rearrangement [27, 94], grasping [95] and SLAM [96]. While these benchmarks ensure reproducibility and fair comparison of different methods, there is always a gap between simulation and reality since it is infeasible to model all details of the real world in simulation. Our benchmark, in contrast, enables fair comparison of different methods and reproducibility of the results in the real world. Additionally, previous benchmarks often operate in a simplified discrete action space [20, 6], even forcing that structure on the real world [2].
@@
Appendix B Metrics
@@
B.1 Simulation Success Metrics
@@
B.2 Real World Success Metrics
@@
Given that the scene we use in the real world is much smaller than the apartments in simulation, we allow the agent to act in the environment for 300 timesteps. The episode is considered to have succeeded if it succeeds in all 4 stages.
Appendix C Simulation Details
C.1 Object Categories Appearing in the Scene Dataset
{tcolorbox} Single seen instance categories: watch, toy swing, toy pineapple, toy food, toy fire truck, toy cactus, sushi mat, stuffed toy, spoon, spectacles, spatula, soap dish, screwdriver, scissors, ramekin, pitcher, mouse pad, monitor stand, milk frother cup, lunch box, laptop cover, lamp, ladle, keychain, hat, handbag, hammer, fork, folder, file sorter, carrying case, candy bar, candle holder, cake pan, c-clamp, butter dish, bundt pan, bath towel, basketball
{tcolorbox} Single instance categories: watering can, video game cartridge, utensil holder cup, toy table, toy sofa, toy sink, toy refrigerator, toy lamp, toy fruits, toy construction set, toy bee, tissue box, squeezer, soap, helmet, electronic cable, doll, dish, can opener, battery charger, backpack
Figure 6: Number of objects across different splits, for both seen categories and unseen categories. We divide objects between categories which appear in training data – seen categories – and those that do not – unseen categories. The goal of Open-Vocabulary Mobile Manipulation is to be able to find and manipulate objects specified by language.
@@ C.1 Object Categories Appearing in the Scene Dataset
C.2 Episode Generation Details
@@
The full set of included receptables in simulation is: bathtub, bed, bench, cabinet, chair, chest_of_drawers, couch, counter, filing_cabinet, hamper, serving cart, shelves, shoe_rack, sink, stand, stool, table, toilet, trunk, wardrobe, & washer_dryer.
C.3 Improved scene visuals
@@
We also benchmark the ObjectNav training speeds of a DDPPO-based RL agent with and without the improved rendering and present the results in 12. We see that the improvement in scene lighting and rendering comes at the cost of only a 3% dip in training FPS (decreasing from around 340 to around 330).
@@
C.4 Action Space Implementation
@@
Continuous. Our continuous action space is implemented as a teleporting agent, where the robot needs to move around by predicting a local waypoint. Our robot’s low level controllers are expected to be able to get the robot to this location, in lieu of simulating full physics for the agent.
@@
Appendix D HomeRobot Implementation Details
@@
  • home_robot: Shared components such as Environment interfaces, controllers, detection and segmentation modules.
@@
  • home_robot_hw: Hardware stack with server processes that runs on the robot, client API that runs on the GPU workstation, and Environments built using the client API.
@@
  • Agents contain all of the necessary code to execute policies. We implement agents which use a mixture of heuristic policies and policies learned on our scene dataset via reinforcement learning.
  • Environments provide common logic; they provide Observations to the Agent, and a function which allows them to apply their action to the (real or simulated) environment.
D.1 Pose Information
@@
D.2 Low-Level Control for Navigation
The Hello Stretch software provides a native interface for controlling the linear and angular velocities of the differential-drive robot base. While we do expose an interface for users to control these velocities directly, it is desireable to have desired short-term goals as a more intuitive action space for policies, and to make them update-able at any instant to allow for replanning.
Thus, we implemented a velocity controller that produces continuous velocity commands that moves the robot to an input goal pose. The controller operates in a heuristic manner: by rotating the robot so that it faces the goal position at all times while moving towards the goal position, and then rotating to reach the goal orientation once goal position is reached. The velocities to induce these motions are inferred with a trapezoidal velocity profile and some conditional checks to prevent it from overshooting the goal.
@@ D.2 Low-Level Control for Navigation
The Fast Marching Method-based motion planning from prior work [2] that we describe in Sec. D.2. It assumes the agent is a cylinder, and therefore is much more limited in where it can navigate than, e.g., a sampling based motion planner like RRT-connect [100] which can take orientation into account. In addition, our semantic mapping requires a list of classes for use with DETIC [26]; instead, it would be good to use a fully open-vocabulary scene representation like CLIP-Fields [11], ConceptFusion [15], or USA-Net [12], which would also improve our motion planning significantly.
D.3 Heuristic Grasping Policy
@@
Numerous powerful grasp generation models have been proposed in the literature, such as GraspNet-1Billion [66], 6-DOF GraspNet [65], and Contact-GraspNet [64]. However, for transparency, reproducibility, and ease of installation, we implement a simple, heuristic grasping policy, which assumes a parallel gripper performing top-down grasps. Heuristic grasp policies appear throughout robotics research (e.g. in TidyBot [62]). In our case, the heuristic policy voxelizes the point cloud, and chooses areas at the top of the object where points exist, surrounded by free space, in order to grasp. Fig. 13 shows the simple grasp policy in action and additional details are presented in Sec. D.3. This policy works well on a wide variety of objects, and we saw comparable performance to the state-of-the-art open-source grasping models we tested [64, 66].
The intuition is to identify areas where the gripper fingers can descend unobstructed into two sides of a physical part of the object, which we do through a simple voxelization scheme. We take the top 10% of points in an object, voxelize at a fixed resolution of 0.5cm, and choose grasps with free voxels (where fingers can go) on either side of occupied voxels. In practice, this achieved a high success rates on a variety of real objects.
@@
We compared this policy to other methods like ContactGraspnet [64], 6-DoF Graspnet [65, 97], and Graspnet 1-Billion [66]. We saw more intermittent failures due to sensor noise using these pretrained methods, even after adapting the grasp offsets to fit to the Hello Robot Stretch’s gripper geometry. In the end, we leave training better grasp policies to future work.
D.4 Heuristic Placement Policy
@@ D.6 Navigation Limitations
Appendix E Reinforcement Learning Baseline
@@
E.1 Action Space
E.1.1 Navigation Skills
FindObject and FindReceptacle are, collectively, navigation skills. For these two skills, we use the discrete action space, as mentioned in Sec. C.4. In our experiments, we found the discrete action space was better at exploration and easier to train.
E.1.2 Manipulation Skills
@@ E.3.1 Navigation Skills
Initial state: The agent is spawned at least 3m away from candidate object or receptacle. It starts in “navigation mode,” with the robot’s head facing forward.
@@
Success condition: The agent should call the discrete stop action when it reaches within 0.5m of a goal view point. The agent should be facing the target: the angle between agent’s heading direction and the ray from robot to center of the closest candidate object should be no more than 15 degrees.
@@
E.3.2 GazeAtObject
The GazeAtObject skill starts near the object, and provides some final refinement steps until the agent is close enough to call a grasp action, i.e. it is in arm’s length of the object and the object is centered and visible. The agent needs to move closer to the object and then adjust its head tilt until the candidate object is close and centered. It makes predictions to move and rotate the head, as well as to center the object and make sure it’s within arm’s length, so that the discrete grasping policy can execute.
The GazeAtObject skill is supposed to start off from locations and help reach a location within arm’s length of a candidate object. This is trained by first initialising the agents close to candidate start receptacles. The agent is then tasked to reach close to the agent and adjust its head tilt such that the candidate object is close and centered in the agent’s camera view. We next provide details on the training setup.
@@ E.3.2 GazeAtObject
Reward: We train the gaze-policy mainly with a dense reward based on distance to goal. Specifically, assuming the distance of the end-effector to the closest candidate goal at time $t$ is $d(t)$ (in metres), the agent receives a reward proportional to $d(t-1)-d(t)$. Further, when the agent reaches with 0.8m, we provide an additional reward for incentivizing the agent to look towards the object.
Let $\theta(t)$ denote the angle (in radians) between the ray from agent’s camera to the object and camera’s normal. Then the reward is given as:
@@
The agent receives an additional positive reward of $2$ once the episode succeeds and receives a negative reward of $-0.5$ for centering its camera towards a wrong object.
E.3.3 PlaceObject
@@
Input observations: Ground truth segmentation of goal receptacles, head depth sensor, joint sensor, sensor indicating if the agent is holding any object, CLIP [14] embedding for the name of object being held.
@@
Actions: Base translation and rotation waypoints, all arm joints (arm extension, arm lift, gripper yaw, pitch and roll), a manipulation mode action that can be invoked only once in an episode to turn the agent’s head towards the arm and rotate the base left by 90 degrees. The agent is not allowed to move its base while in manipulation mode.
@@
Reward: The agent receives a positive sparse reward of $5$ when it releases the object and the object comes in contact with a target receptacle. Additionaly, we provide a positive reward of $1$ for each step the object stays in contact with the target receptacle. It receives a negative reward of $-1$ if the agent releases the object but the object does not come in contact with the receptacle.
Appendix F Additional Analysis

### v2

@@ 6 Conclusions and Future Work
We proposed a combined simulation and real-world benchmark to enable progress on the important problem of Open-Vocabulary Mobile Manipulation. We ran extensive experiments showing promising simulation and real-world results from two baselines: a heuristic baseline based on a state-of-the-art motion planner [2] and a reinforcement learning baseline trained with DDPPO [64]. In the future, we hope to improve the complexity of the problem space, adding more complex natural language and multi-step commands, and provide end-to-end baselines instead of modular policies.
@@ 7 Acknowledgements
The Georgia Tech effort was supported in part by ONR YIP and ARO PECASE. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of the U.S. Government, or any sponsor.
@@ Appendix A Extended Related Work
Datasets. In addition to the environments referenced in Table 3, offline datasets including robot interactions with scenes have been used widely to train models. These datasets are typically obtained using robots alone (e.g., [73, 74]), by teleoperation (e.g., [75, 76]) or human-robot demonstration (e.g., [77]). Previous works such as [78] aim to collect large-scale datasets while works such as [79] consider scaling across multiple embodiments. [80] take a step further by collecting robot data in unstructured environments. Unlike these works, we do not limit our users to a specific dataset. Instead, we provide a simulator with various scenes that can generate large-scale consistent data for training. Also, note that we test the models in unseen environments, while most of the mentioned works use the same environment for training and testing.
Simulation benchmarks. The embodied AI community has provided various benchmarks in simulation platforms for tasks such as navigation [1, 81, 82, 83, 84], object manipulation [85, 35, 86, 87], instruction following [6, 88, 89, 90], room rearrangement [28, 91], grasping [92] and SLAM [93]. While these benchmarks ensure reproducibility and fair comparison of different methods, there is always a gap between simulation and reality since it is infeasible to model all the details of the real world in simulation. Our benchmark, in contrast, enables fair comparison of different methods and reproducibility of the results in the real world. Additionally, previous benchmarks often operate in a simplified discrete action space [20, 6], even forcing that structure on the real world [2].
@@
Exploration of unseen environments. Various papers have looked at the problem of object search in different home environments. Gervet et al. [2] use a mixture of heuristic/model-based planning and reinforcement learning to achieve strong results in a variety of real-world environments; importantly their planning-based methods perform competitively to the best learning-based methods, and much better than end-to-end reinforcement learning on real tasks. Also promising is graph-based exploration. Kurenkov et al. [94] propose hierarchical mechanical search, based on a 3d scene graph representation, for exploring environments; although unlike in our Open-Vocabulary Mobile Manipulation task, they assume such a scene graph exists. Similarly, SayPlan [95] performs a search over a complex scene graph using a large language model; however, this approach also does not look into iteratively constructing this scene graph on new scenes. While there is active work in iteratively exploring and building scene graphs and other hierarchical representations (e.g. [96]), there are not yet strongly established methods in this space.
Appendix B Limitations
Due to simulation limitations, we don’t physically simulate grasping in the current version of the benchmark, which is why we provide a separate policy for this in the real world. Grasping is a well-studied problem [97, 98, 99], but simulations that train useful real-world grasp systems require special consideration. We also currently consider full natural language queries out-of-scope. Finally, we did not evaluate many motion planners (see Sec. E.2), or performed task-and-motion-planning with replanning, as would be ideal for a long horizon task [100].
Appendix C Metrics
@@
C.1 Simulation Success Metrics
@@
Better pick success condition. We plan to use a more realistic grasping condition in simulation. We try replacing the magic snap in simulation with a stricter condition that requires the agent to move its arm near the object without colliding with the scene or other objects. Additionally, we tested a baseline (Figure 5) that performs top-down grasps resembling our real-world grasping policy, and resorting to side-ways grasps when the object is farther. While this baseline succeeds in reaching the object starting from an object viewpoint 79% of the time, it does so without colliding only 47% of the time.
Figure 5: A few success and failure cases for our simple grasping policy under the new grasp success condition that requires the agent’s arm to reach near the object without colliding. The agent resorts to sideways grasps when the object can’t be reached via a top-down grasp that bends the gripper. Most grasping failures are because of the collisions with the scene.
C.2 Real World Success Metrics
@@
Given that the scene we use in the real world is much smaller than the apartments in the simulation, we allow the agent to act in the environment for 300 timesteps. The episode is considered to have succeeded if it succeeds in all 4 stages.
Appendix D Simulation Details
D.1 Object Categories Appearing in the Scene Dataset
Figure 6: Number of objects across different splits, for both seen categories and unseen categories. We divide objects between categories that appear in training data – seen categories – and those that do not – unseen categories. The goal of Open-Vocabulary Mobile Manipulation is to be able to find and manipulate objects specified by language.
@@ D.1 Object Categories Appearing in the Scene Dataset
D.2 Episode Generation Details
@@
The full set of included receptacles in simulation is: bathtub, bed, bench, cabinet, chair, chest_of_drawers, couch, counter, filing_cabinet, hamper, serving cart, shelves, shoe_rack, sink, stand, stool, table, toilet, trunk, wardrobe, & washer_dryer.
D.3 Diversity in Receptacle Instances
The instances within each receptacle category exhibit substantial variability. Figure 11 shows a few different receptacles from our dataset belonging to the "table" category.
Figure 11: The variation in instances belonging to the "table" category in our dataset.
D.4 Scene Clutter Complexity
Our procedural placement of target and distractor objects creates diverse and interesting scenarios that require reasoning over which direction to approach receptacles, stable placement in clutter, open vocabulary object detection under occlusion which makes the task quite challenging. Figure 12 shows a few examples of clutter surrounding target objects in our scenes.
Figure 12: A few examples of clutter surrounding the target object in our simulation settings.
D.5 Improved scene visuals
@@
We also benchmark the ObjectNav training speeds of a DDPPO-based RL agent with and without the improved rendering and present the results in 14. We see that the improvement in scene lighting and rendering comes at the cost of only a 3% dip in training FPS (decreasing from around 340 to around 330).
@@
D.6 Action Space Implementation
@@
Continuous. Our continuous action space is implemented as a teleporting agent, where the robot needs to move around by predicting a local waypoint. Our robot’s low-level controllers are expected to be able to get the robot to this location, in lieu of simulating full physics for the agent.
@@
Appendix E HomeRobot Implementation Details
@@
  • home_robot: Shared components such as Environment interfaces, controllers, detection, and segmentation modules.
@@
  • home_robot_hw: Hardware stack with server processes that run on the robot, client API that runs on the GPU workstation, and Environments built using the client API.
@@
  • Agents contain all of the necessary code to execute policies. We implement agents that use a mixture of heuristic policies and policies learned on our scene dataset via reinforcement learning.
  • Environments provide common logic; they provide Observations to the Agent and a function that allows them to apply their action to the (real or simulated) environment.
E.1 Pose Information
@@
E.2 Low-Level Control for Navigation
The Hello Stretch software provides a native interface for controlling the linear and angular velocities of the differential-drive robot base. While we do expose an interface for users to control these velocities directly, it is desirable to have desired short-term goals as a more intuitive action space for policies, and to make them updateable at any instant to allow for replanning.
Thus, we implemented a velocity controller that produces continuous velocity commands that move the robot to an input goal pose. The controller operates in a heuristic manner: by rotating the robot so that it faces the goal position at all times while moving towards the goal position, and then rotating to reach the goal orientation once the goal position is reached. The velocities to induce these motions are inferred with a trapezoidal velocity profile and some conditional checks to prevent it from overshooting the goal.
@@ E.2 Low-Level Control for Navigation
The Fast Marching Method-based motion planning from prior work [2] that we describe in Sec. E.2. It assumes the agent is a cylinder, and therefore is much more limited in where it can navigate than, e.g., a sampling-based motion planner like RRT-connect [104] which can take orientation into account. In addition, our semantic mapping requires a list of classes for use with DETIC [27]; instead, it would be good to use a fully open-vocabulary scene representation like CLIP-Fields [11], ConceptFusion [15], or USA-Net [12], which would also improve our motion planning significantly.
E.3 Heuristic Grasping Policy
@@
Numerous powerful grasp generation models have been proposed in the literature, such as GraspNet-1Billion [99], 6-DOF GraspNet [98], and Contact-GraspNet [97]. However, for transparency, reproducibility, and ease of installation, we implement a simple, heuristic grasping policy, which assumes a parallel gripper performing top-down grasps. Heuristic grasp policies appear throughout robotics research (e.g. in TidyBot [62]). In our case, the heuristic policy voxelizes the point cloud, and chooses areas at the top of the object where points exist, surrounded by free space, in order to grasp. Fig. 15 shows the simple grasp policy in action and additional details are presented in Sec. E.3. This policy works well on a wide variety of objects, and we saw comparable performance to the state-of-the-art open-source grasping models we tested [97, 99].
The intuition is to identify areas where the gripper fingers can descend unobstructed into two sides of a physical part of the object, which we do through a simple voxelization scheme. We take the top 10% of points in an object, voxelize at a fixed resolution of 0.5cm, and choose grasps with free voxels (where fingers can go) on either side of occupied voxels. In practice, this achieved a high success rate on a variety of real objects.
@@
We compared this policy to other methods like ContactGraspnet [97], 6-DoF Graspnet [98, 101], and Graspnet 1-Billion [99]. We saw more intermittent failures due to sensor noise using these pretrained methods, even after adapting the grasp offsets to fit to the Hello Robot Stretch’s gripper geometry. In the end, we leave training to better grasp policies for future work.
E.4 Heuristic Placement Policy
@@ E.6 Navigation Limitations
Appendix F Reinforcement Learning Baseline
@@
F.1 Action Space
F.1.1 Navigation Skills
FindObject and FindReceptacle are, collectively, navigation skills. For these two skills, we use the discrete action space, as mentioned in Sec. D.6. In our experiments, we found the discrete action space was better at exploration and easier to train.
F.1.2 Manipulation Skills
@@ F.3.1 Navigation Skills
Initial state: The agent is spawned at least 3m away from the candidate object or receptacle. It starts in “navigation mode,” with the robot’s head facing forward.
@@
Success condition: The agent should call the discrete stop action when it reaches within 0.5m of a goal viewpoint. The agent should be facing the target: the angle between the agent’s heading direction and the ray from the robot to the center of the closest candidate object should be no more than 15 degrees.
@@
F.3.2 GazeAtObject
The GazeAtObject skill starts near the object and provides some final refinement steps until the agent is close enough to call a grasp action, i.e. it is in arm’s length of the object and the object is centered and visible. The agent needs to move closer to the object and then adjust its head tilt until the candidate object is close and centered. It makes predictions to move and rotate the head, as well as to center the object and make sure it’s within arm’s length so that the discrete grasping policy can execute.
The GazeAtObject skill is supposed to start off from locations and help reach a location within arm’s length of a candidate object. This is trained by first initializing the agents close to candidate start receptacles. The agent is then tasked to reach close to the agent and adjust its head tilt such that the candidate object is close and centered in the agent’s camera view. We next provide details on the training setup.
@@ F.3.2 GazeAtObject
Reward: We train the gaze policy mainly with a dense reward based on the distance to the goal. Specifically, assuming the distance of the end-effector to the closest candidate goal at time $t$ is $d(t)$ (in meters), the agent receives a reward proportional to $d(t-1)-d(t)$. Further, when the agent reaches 0.8m, we provide an additional reward for incentivizing the agent to look toward the object.
Let $\theta(t)$ denote the angle (in radians) between the ray from the agent’s camera to the object and the camera’s normal. Then the reward is given as:
@@
The agent receives an additional positive reward of $2$ once the episode succeeds and receives a negative reward of $-0.5$ for centering its camera towards the wrong object.
F.3.3 PlaceObject
@@
Input observations: Ground truth segmentation of goal receptacles, head depth sensor, joint sensor, sensor indicating if the agent is holding any object, CLIP [14] embedding for the name of the object being held.
@@
Actions: Base translation and rotation waypoints, all arm joints (arm extension, arm lift, gripper yaw, pitch, and roll), a manipulation mode action that can be invoked only once in an episode to turn the agent’s head towards the arm and rotate the base left by 90 degrees. The agent is not allowed to move its base while in manipulation mode.
@@
Reward: The agent receives a positive sparse reward of $5$ when it releases the object and the object comes in contact with a target receptacle. Additionally, we provide a positive reward of $1$ for each step the object stays in contact with the target receptacle. It receives a negative reward of $-1$ if the agent releases the object but the object does not come in contact with the receptacle.
F.4 ConceptFusion
In the main paper, we introduced two key approaches based on end-to-end reinforcement learning and a heuristic baseline. Both methods are dependent on the detection results generated by a readily available open vocabulary object detector  [27]. Notably, these 2D detection models do not leverage information from prior time steps to inform their detection decisions.
In order to address these limitations, we explored the application of ConceptFusion [15], an open-set scene representation technique. ConceptFusion harnesses foundation models like CLIP [14], DINO [109], and others to construct 3D maps from multiple images. For our experimentation, we employed the open-source implementation of ConceptFusion, which utilizes the Segment Anything Model (SAM) [110] for object segmentation in RGB images and CLIP for feature extraction from each segmentation mask. It’s important to note that our experiments were conducted in a simulated environment, obviating the need for GradSLAM [111], as we had access to ground truth depth maps and pose information to support our map construction efforts.
During our initial experimentation, we observed that ConceptFusion demanded significant computational resources and memory, with processing times reaching up to 5 seconds per frame for map construction. Remarkably, it’s worth noting that the authors of ConceptFusion have recently published a new paper titled "ConceptGraphs: Open-Vocabulary 3D Scene Graphs for Perception and Planning," [112] which addresses some of the computational challenges we encountered. However, we leave the exploration of ConceptGraphs as a potential avenue for future research.
Appendix G Additional Analysis